OpenStudy (anonymous):

8 inches of heavy wet snow are equivalent to an inch of rain, ... estimate water content in 13 inches of heavy wet snow

OpenStudy (anonymous):

8in of snow= 1 inch of rain. 1/8=? 13/8=?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

@abhrag321

OpenStudy (anonymous):

i believe find out an inch of heavy wet snow is equal to _______ inches of rain. then multiply both sides by 13.

OpenStudy (anonymous):

you are right @reason

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

calculator????

OpenStudy (anonymous):

lol im so lost

OpenStudy (anonymous):

1 iphone = 12 Droid phones. how many droid phones are equal to 3 iPhones? lmao

OpenStudy (anonymous):

13/8=13(25)/8(25)=325/200=(325/2)/100=162.5/100=1.625